1) Saxbury Corporation's relevant range of activity is 3,000 units to 7,000 units. When it produces and sells 5,000 units, its average costs per unit are as follows:

Average Cost per Unit

Direct materials

$5.30

Direct labor

$3.65

Variable manufacturing overhead

$1.50

Fixed manufacturing overhead

$3.90

Fixed selling expense

$0.75

Fixed administrative expense

$0.60

Sales commissions

$0.50

Variable administrative expense

$0.50

Required:

i. If 4,000 units are produced, what is the total amount of direct manufacturing cost incurred?

j. If 4,000 units are produced, what is the total amount of indirect manufacturing cost incurred?

k. What incremental manufacturing cost will the company incur if it increases production from 5,000 to 5,001 units?
Answer #1

1.Direct manufacturing cost = Direct Material + Direct labor

= (5.30+3.65)*4000

= $35,800

j.Indirect manufacturing cost = Variable Manufacturing Overheads + Fixed Manufacturing Overhead’

= 1.50*4000+3.90*5000

= $25,500

k.Incremental manufacturing cost = variable manufacturing cost as fixed costs do not change with change in output within the relevant range.

Hence, incremental manufacturing cost = 5.30+3.65+1.50

= $10.45